OpenStudy (anonymous):

Need help w/ a Maclaurin series: f(x)=(1-x)^(-2) I know how to do the derivation and write out the series. I mostly need help w/ understanding and writing the summation.

OpenStudy (anonymous):

\[f(x)=\frac{1}{(1-x)^2}\] well, the geometric series looks like \[\sum_{0}^{\infty} x^n=\frac{1}{1-x}\] the derivative of tfe right hand side is \[\frac{1}{(1-x)^2}=\sum_{1}^{\infty} nx^{n-1}\] and there you go, just take derivative of summation with respect to x, and change the lower index by +1.

OpenStudy (anonymous):

Feel free to ask if in doubt of anything.

OpenStudy (anonymous):

Okay, so it is like doing a power series? And so the summation above would be the final summation? Or do I still need to manipulate it?

OpenStudy (anonymous):

And thank you btw for answering my question! :D

OpenStudy (anonymous):

Oh wait, I think the only thing I am confused about is where the "n" comes from (before the x^(n-1)) in the last summation above.

OpenStudy (anonymous):

So "n" is the last power, so for example \[(\frac{1}{2})^{0}+(\frac{1}{2})^{1}+(\frac{1}{2})^{2}+(\frac{1}{2})^{3}+(\frac{1}{2})^{4}+....+(\frac{1}{2})^{n-2}+(\frac{1}{2})^{n-1}+(\frac{1}{2})^{n}=\sum_{k=0}^{n}(\frac{1}{2})^k\] The "k" is just the powers, so 0,1,2,3,4.... all the way up to the last one which we call "n".So n is just the last power, so in our series the "n" goes to infinity.

OpenStudy (anonymous):

I get it now :) Thanks again!